In a ISOSCELES TRIANGLE
m∠T = m∠U = (2x)°,
find m∠V (in degrees) as an expression containing variable x
Yes, this answer will have a variable.

Answer by ikleyn(52788)

.

    m∠V (in degrees) = 180 - 2x - 2x = 180 - 4x.    ANSWER

Solved.